CHALLENGE ACTIVITY 5.7.1: Matrix representation with respect to nonstandard bases. Jump to level 1 1 2 Let T : R3 + R2 be defLinear
Answer #1

Solution - T: IR→IR? w T.1 ) 6-$4,- (3]..:]7:13 = { vin ( 2 , 02 (373 [31] [ *114-57-(2/3) 777B) 3 - 12 + 7 - 12 -5 = - 7A*-*-*-*-* = - 1 + 3 3 - S A - Q - ) we find eqro + 3xeqno 7A + SA- = و9+ ا ) 3 = 8 A د ا = A تمام عام را : 13 د لا Hence (4۔